OpenStudy (anonymous):

I have three questions I need help on, if anyone is willing to help me with. I would greatly appreciate it. Will Medal. Please and thank you.

OpenStudy (anonymous):

1. What is the altitude of the cone?

OpenStudy (anonymous):

2. What is the volume of the cone? I used google calculator for this one and the last one and the numbers do not match. I'm getting 50.27 for this one. but here are the answers. a. 60 b. 20 c. 16 d. 54

OpenStudy (anonymous):

3. What is the surface area of the cone? I'm getting 301.59 what am I doing wrong?

OpenStudy (anonymous):

answers for the last one, I forgot to list. a. 96 b.20 c.36 d. 60

Awolflover1 (awolflover1):

1. obviously, CE is altitude of that cone,

Awolflover1 (awolflover1):

The distance from the apex (vertex) of a cone to the base is the altitude. It is the shortest (perpendicular distance) line segment between the apex of a cone and the (possibly extended) base. Altitude also refers to the length of this segment. So, that would be (as already given without explanation :) ) segment CE.

OpenStudy (anonymous):

Thank you @Awolflover1! :)

OpenStudy (anonymous):

surface area of the cone is pi* radius * slant height + pi* square of radius.

Awolflover1 (awolflover1):

2. I think its A Because i times 5x3x4=60

OpenStudy (anonymous):

@cooldude17 says it's 16 pi. Uhoh, so is it 60 or 16?

Awolflover1 (awolflover1):

Hmm

OpenStudy (anonymous):

its 1/3 pi * 4*4*3 = 16 pi cm cube

OpenStudy (anonymous):

calculate the surface area.

OpenStudy (anonymous):

Would I do it the same way you calculated the volume or no?

OpenStudy (anonymous):

no surface area has got a different formula

OpenStudy (anonymous):

Would it be 36 pi cm3?

OpenStudy (anonymous):

cm2*

OpenStudy (anonymous):

pi*r*l = pi*6*8 + pi*6*6 = 36pi+48pi = 74pi cm square.

OpenStudy (anonymous):

74 isn't one of the choices?

OpenStudy (anonymous):

sorry, it definitely has to be 84 pi cm 2

OpenStudy (anonymous):

84 is not a choice either. @cooldude17

OpenStudy (anonymous):

what do you think @Awolflover1

OpenStudy (anonymous):

the choices are. 96 20 36 60

OpenStudy (anonymous):

so most rational choice comes to be 36

OpenStudy (anonymous):

alright well thank you!

OpenStudy (anonymous):

when u talk about surface area.. iits curved surface area + surface area of base.

OpenStudy (anonymous):

Well for the three questions you guys helped me with I still got 9 out of 11 questions correct.

OpenStudy (anonymous):

But it don't tell me which ones I got incorrect.

OpenStudy (anonymous):

hey u got more questions?